Consider a quadratic function that has a minimum value of -6, and x-intercepts at (-6,0) and (0,0).
A)Determine the equation of the quadratic function. Show or explain your work.

....please help me! I really want to understand how to do this. I'm not sure how and what the answer would be.

Answer by stanbon(75887)

Consider a quadratic function that has a minimum value of -6, and x-intercepts at (-6,0) and (0,0)
---------
Plot the x-intercepts
Since the minimum y-value is -6
the vertex must be at (-3,-6)
=============================
y = a(x-0)(x-(-6))
Solve for "a":
-6 = a(-3)(-3+6)
----
-6 = a(-3)(3)
a = 6/9 = 2/3
----
y = (2/3)x(x+6)
-----
y = (2/3)(x^2+6x)
-
y = (2/3)x^2 + 4x
========================
Cheers,
Stan H.
========================